# Quantum Complexity of Nonlocal Field Theories

### v2

@@ 3.1 Volume complexity for the anisotropic case
Where, length scale of noncommutativity is $a=\lambda^{1/4}\sqrt{\theta}$, $\lambda$ is (dimensionless) t’Hooft coupling, $\sqrt{\theta}$ is the noncommutativity parameter and $B_{xy}$ is the only non-zero component of the NS-NS form background. The volume complexity can be found by taking a space like hypersurface specified by the embedding $t=t(u)$ in the string frame as,
@@ 3.2 Subregion Volume complexity (CV) for the anisotropic case
To compute subregion complexity we focus our attention on the portion of the maximal volume slice which is contained within the Ryu-Takayanagi (RT) surface, and which is also homologous to the boundary subregion of interest. We take the bondary subregion to be the infinitely long strip, $\frac{l}{2}<x<\frac{l}{2},\ -\frac{W}{2}<y,z<\frac{W}{2}$ in the limit $W\to\infty$ with $l$ fixed. In this case the pullback metric on the bounding codimension 2 (RT) surface is
@@ 3.2.1 Near AdS/Weakly noncommutative regime
$C_{V}$ $=\frac{W^{2}}{G_{N}^{(5)}}\left(\frac{lu_{b}^{3}}{3}-\frac{4\pi^{3/2}\Gamma\left(\frac{2}{3}\right)}{9l^{2}\Gamma\left(\frac{1}{6}\right)}+a^{4}\left(\frac{128\pi^{9/2}\Gamma\left(\frac{2}{3}\right)^{9}}{9l^{9}u_{b}^{3}\Gamma\left(\frac{1}{6}\right)^{9}}-\frac{8\pi^{3/2}u_{b}^{3}\Gamma\left(\frac{2}{3}\right)^{3}}{9l^{3}\Gamma\left(\frac{1}{6}\right)^{3}}\right)+O(a^{8})\right)$ (84)
@@ 3.2.2 Far AdS/extreme noncommutative regime
$C_{V}$ $=\frac{1}{G_{N}^{(5)}}\left(\frac{W^{2}lu_{b}^{3}}{3}+\frac{W^{2}l^{4}}{a^{6}}\frac{\sqrt{3}\Gamma\left(\frac{5}{6}\right)^{6}}{\sqrt[3]{2}\pi^{3}\Gamma\left(\frac{4}{3}\right)^{3}}-\frac{W^{2}l^{3}u_{b}}{a^{4}}\frac{\Gamma\left(-\frac{1}{6}\right)^{4}}{32\sqrt[3]{2}\sqrt{3}\pi^{3}\Gamma\left(\frac{1}{3}\right)}\right)$ (86)
@@ 3.3 Action complexity (CA) for the anisotropic case (10 D)
    We also need to incorporate the appropriate topological boundary term Kurlyand:2022vzv ; Apolo:2025wcl to get correct action leading to the gravitational background with a negative cosmological constant. Topological boundary term is required to correctly account for the self dual five-form, $F_{5}$ field and can be motivated by decomposing the $F_{5}$ into electric and magnetic parts in the following manner
@@
    $S_{\partial\mathcal{V}}$ $=\sum_{j=\pm}\frac{2}{16\pi G_{N}}\int d^{3}x\int d\lambda_{j}\,\sqrt{\gamma_{j}}\,\kappa_{j}.$ (94)
    Here, the affine parameter $\lambda$ along the null generators increases in the future direction, and $\gamma$ represents the induced metric on constant-$\lambda$ slices. On the section of constant $\lambda_{\pm}=\mp u$,
@@
    where the ++ and - indices denote the upper and lower null boundaries, respectively. The affine parameters are chosen such that $\lambda_{\pm}=\mp u$, leading to the following null normal vectors
@@
    In the perturbative limit where $a\ll 1/r_{b}$, this expression receives vanishing contribution up to $\mathcal{O}(a^{4})$ in non-commutative coupling. This is consistent with the the expected AdS result.
    In contrast, the strongly non-commutative regime (where $a\gg 1/r_{b}$) contribution is given by
    $S_{\partial\mathcal{V}}^{NC}$ $=\frac{L^{3}u_{b}^{3}}{16\pi G_{N}}\left(-\frac{4}{3}\alpha\right)+O\left(\frac{1}{a^{3}}\right).$ (96)
  • LMPS counterterm prescription: As noted in Lehner:2016vdi , the gravitational action depends on the choice of the affine parameter $\lambda$ on null boundaries. To preserve reparametrization invariance, the following counterterm must be included for each null boundary
    $\sum_{\pm}\Delta S_{i}$ $=-2\int d{A_{i}}\,d\lambda_{i}\,{\Theta_{i}}\ln|l_{ct}{\Theta_{i}}|,\hskip 85.35826pt\left({\Theta_{i}}\equiv\frac{1}{\sqrt{{\gamma}}}\frac{\partial\sqrt{{\gamma}}}{\partial\lambda_{i}}\right).$
    Here, $\gamma$} denotes the induced metric on constant-$\lambda$ cross-sections, which form codimension-2 surfaces. On slices with constant $\lambda_{\pm}=\mp r$, the integration measure is as derived earlier ($\sqrt{\gamma}=u^{3}h(u)$). After accounting for sign conventions ($\Theta_{+}=-\Theta_{-}$), and to the causal structure, the total contribution from both surfaces turns out to be
    $\Delta S$ $=\frac{4}{16\pi G_{N}}\int d{A}\,\int_{0}^{\infty}d\lambda_{-}\,{\Theta}_{-}\ln|{l_{ct}\Theta}_{-}|,$ (97)
@@
$\mathcal{C}$ $=\frac{L^{3}u_{b}^{3}}{16\pi^{2}G_{N}}\left(4\ln 3\right)+O(a^{4}),$ (103)
where, consistency with the pre-established AdS result Reynolds:2016rvl requires us to fix $\alpha$ to be equal to $l_{ct}$.
@@
$\mathcal{C}$ $={\frac{L^{3}u_{b}^{3}}{16\pi^{2}G_{N}}\left(-2\ln a+\frac{16}{3}\right)+O\left({a^{-1}}\right).}$ (104)
The arguments of logarithm have been rendered dimensionless by rewriting them in the units of AdS radius. Moreover, $l_{ct}$ has been chosen to be one in the units of AdS radius without tempering the known universal result for the local field theory dual to AdS bulk. It can be seen that for the anisotropic case of noncommutative field theory, the action complexity is positive only in the range $\ln a<\frac{8}{3}$. Conversely positivity of action complexity imposes an upper bound on the noncommutativity parameter $a<\exp{\left(8/3\right)}$. This feature of action complexity is parallels with the action complexity of the DDSYM theory. However, unlike the DDSYM theory, but instead just like the local SYM field theory, the UV divergence scales with volume $L^{3}$! We defer our speculations about this counterintuitive result till Sec. 5.
@@ 4 Noncommutative SYM theory: Isotropic case
$ds_{E}^{2}=l^{2}\,{r^{2}}\,\Phi\,\left(\Phi^{-2}\left(dx_{0}^{2}+dx_{1}^{2}+dx_{2}^{2}+dx_{3}^{2}\right)+r^{-4}dr^{2}+r^{-2}d\Omega^{2}_{5}\right)$
$\Phi$ $=(1+a^{4}r^{4})^{1/2},\quad F_{0123r}=i4l^{4}r^{3}\Phi^{-4},$
$B_{01}=B$ $_{23}=\sqrt{g_{s}}a^{2}r^{4}\Phi^{-2},\quad A_{01}=A_{23}=-i\frac{a^{2}l^{2}}{\sqrt{g_{s}}}r^{4}\Phi^{-2},\quad\chi=i\frac{a^{4}r^{4}}{g_{s}}$ (105)
@@
$ds^{2}$ $=l^{2}r^{-2}\Phi^{8/3}dr^{2}+l^{2}r^{2}\Phi^{2/3}(-dt^{2}+dx_{1}^{2}+dx_{2}^{2}+dx_{3}^{2})$ (106)
where $\Phi=(1+a^{4}r^{4})^{1/2}$. We will work in spherical polar coordinates ($\rho,\theta,\phi$) instead of the cartesian $x_{1},x_{2},x_{3}$ and follow a similar procedure as above to compute the volume complexity.
@@ 4.1 Volume complexity of the isotropic NCSYM
$V$ $=\int dr\ d\rho\ d\theta\ d\phi\ l^{4}r^{4}\Phi^{4/3}\rho^{2}\sin\theta\sqrt{\frac{\Phi^{2}}{r^{4}}-t^{\prime 2}(r)}=\frac{4\pi l^{4}R^{3}}{3}\int_{0}^{1/\epsilon}dr\ r^{2}\Phi^{7/3}$ (107)
@@ 4.2 Action complexity (CA) for the isotropic case (5D)
$\mathcal{R}$ $=-\frac{20\left(3a^{8}r^{8}+10a^{4}r^{4}+3\right)}{3\left(a^{4}r^{4}+1\right)^{10/3}}.$
@@
$dt^{2}$ $=\frac{dr^{2}}{r^{4}}\Phi^{2},$
$\int_{t_{-}}^{t_{+}}dt$ $=2\int^{\Lambda}_{r}dr^{\prime}\,\frac{\Phi(r^{\prime})}{r^{\prime 2}}.$
@@
  • Computation of bulk action: The bulk action is evaluated over WdW patch with integration measure $\sqrt{-g}=r^{3}\Phi^{\frac{8}{3}}$ and gives
@@
    $=\frac{1}{16\pi G_{N}}\int d^{3}x\int_{0}^{\Lambda}dr\int_{t_{-}}^{t_{+}}dt\,\,r^{3}\left(a^{4}r^{4}+1\right)^{4/3}\left(\frac{8\left(2-5\left(a^{4}r^{4}+1\right)^{2}\right)}{3\left(a^{4}r^{4}+1\right)^{10/3}}\right).$
    Working perturbatively, prior to capturing the non-commutative answer, we always first verify the AdS side of the story. Evaluating the bulk action in the perturbative limit when, the UV cutoff scale dominates the non-commuative coupling i.e. $a\ll 1/r_{b}$, reproduces the expected result for the known case of pure AdS
    $S_{bulk}$ $=\frac{V_{3}r_{b}^{3}}{16\pi G_{N}}\left(-\frac{4}{3}+O(a^{4})\right).$ (111)
    Once, we have the correct match with the expected scaling of the known AdS result, we next evaluate the non-commutative answer in the perturbative limit when the non-commutative coupling dominates the UV cutoff scale or in other words, $a\gg 1/r_{b}$. We obtain,
    $S_{bulk}$ $=\frac{V_{3}}{16\pi G_{N}}\left(-\frac{2}{3}a^{2}r_{b}^{5}+O(a^{-2})\right).$ (112)
@@
    $S_{boundary}$ $=\sum_{j=\pm}\frac{2}{16\pi G_{N}}\int d^{3}x\int d\lambda_{j}\,\sqrt{\gamma_{j}}\,\kappa_{j}.$ (113)
    Here, the parameters $\lambda$ on null generators are chosen to be increasing in the future direction, $\gamma$ is the induced metric on the cross-sections of constant $\lambda$. Moreover, $n^{\mu}\nabla_{\mu}n^{\nu}=\kappa n^{\nu}$, where, $n$ is the null normal to the surface. Boundaries of WdW are given by Null surfaces $\mathcal{N}_{\pm}$
    $(t-T)$ $=\pm\int^{\Lambda}_{r}dr^{\prime}\,\frac{\Phi(r^{\prime})}{r^{\prime 2}}.$
@@
    $t$ $=T\pm\int^{\Lambda}_{r}dr^{\prime}\,\frac{\Phi(r^{\prime})}{r^{\prime 2}},$
    where, +/+/- labels the upper/lower null boundaries, respectively. The parameters on null generators are chosen to be increasing in the future direction and are taken to be $\lambda_{\pm}=\mp r$. With this parametrization, the null normal vectors takes the following form
    $n_{\pm}^{\mu}=\left(\frac{\alpha\Phi(r)}{r^{2}},\mp\alpha,\vec{0}\right),$
@@
    $\kappa_{\pm}$ $=\mp\frac{10\alpha a^{4}r^{3}}{3\left(a^{4}r^{4}+1\right)}.$
@@
    $S_{boundary}$ $=\frac{1}{16\pi G_{N}}\frac{40\alpha V_{3}a^{4}}{3}\int_{0}^{r_{b}}dr\,\frac{r^{6}}{\left(a^{4}r^{4}+1\right)},$
    where, on the sections of constant $\lambda_{\pm}$, the induced metric is
    $d\gamma^{2}$ $=r^{2}(dx^{2}+dy^{2}+dz^{2})\Phi^{2/3},$
    giving the integral measure of $\sqrt{\gamma}=r^{3}\Phi(r)$. Working perturbatively, we first verify the AdS side of the story by evaluating the bulk action in the perturbative limit when, $a\ll 1/r_{b}$
    $S_{boundary}$ $=\frac{V_{3}}{16\pi G_{N}}a^{4}\alpha r_{b}^{7}\left(\frac{40}{21}+O(a^{4})\right),$ (114)
    this correctly reproduces the expected result for the known case of pure AdS to the zeroth order in coupling $a$. Now looking ahead and working out the corresponding answer when the non-commutative coupling dominates the UV cutoff scale $a\gg 1/r_{b}$. We obtain,
    $S_{boundary}$ $=\frac{V_{3}}{16\pi G_{N}}\left(\frac{40}{9}\alpha r_{b}^{3}+O\left(\frac{1}{a^{4}}\right)\right).$ (115)
@@
  • LMPS counter-term prescription It has been pointed out in Lehner:2016vdi that the gravitational action depends upon the choice of the parametrization $\lambda$ on the null boundary components. To restore the coordinate independence of the gravitational action, it is required to include the following counterterm component in the action, for each of the null boundaries.
    $\Delta S$ $=\sum_{j=\pm}\frac{-2}{16\pi G_{N}}\int d{A}_{j}\,d\lambda_{j}\,{\Theta}_{j}\ln|l_{ct}{\Theta}_{j}|,\hskip 85.35826pt\text{where,}\,\,{\Theta}_{\pm}\equiv\frac{1}{\sqrt{{\gamma}}}\frac{\partial\sqrt{{\gamma}}}{\partial\lambda_{\pm}}.$ (116)
    Here, $\gamma$} is the induced metric on the constant $\lambda$ sections. Since, metric is defined on the null section, it is a codim-2 surface. And on the section of constant $\lambda_{\pm}=\mp r$, the integration measure is the same as obtained earlier.
@@
    $\text{ }\,\,\Theta$ $=\frac{5a^{4}r^{4}+3}{r\left(a^{4}r^{4}+1\right)}.$
@@
    $\Delta S$ $=\frac{4V_{3}}{16\pi G_{N}}\,\int^{\infty}_{0}dr\,\sqrt{\gamma}\,{\Theta}\ln|l_{ct}{\Theta}|,$
    $=\frac{4V_{3}}{16\pi G_{N}}\,\int_{0}^{r_{b}}dr\,r^{2}\frac{5a^{4}r^{4}+3}{\left(a^{4}r^{4}+1\right)^{1/2}}\ln\left|\frac{l_{ct}(5a^{4}r^{4}+3)}{r\left(a^{4}r^{4}+1\right)}\right|.$
@@
    $\Delta S$ $=\frac{V_{3}r_{b}^{3}}{16\pi G_{N}}\left[\frac{4}{3}+4\ln\left(\frac{3l_{ct}}{r_{b}}\right)\right]+O(a^{4}).$
@@
    $\Delta S^{NC}$ $=\frac{V_{3}}{16\pi G_{N}}\left(2a^{2}r_{b}^{5}\left(-\ln\left(\frac{r_{b}}{l_{ct}}\right)+\ln\left(5\right)+1\right)\right)+O(a^{-2}).$ (117)
@@
    $S_{\mathcal{X}}$ $=\frac{-2}{16\pi G_{N}}\int d^{3}x\,\sqrt{\gamma}\ln\left|\frac{n_{+}\cdot n_{-}}{2}\right|\Bigg|_{r=r_{b}}.$
@@
    $d\gamma^{2}$ $=r^{2}(dx^{2}+dy^{2}+dz^{2})\Phi^{2/3},$
    With the integral measure $\sqrt{\gamma}=r^{3}\Phi(r)$. Given that the various normals have the following components,
    $n^{\mu}_{\pm}$ $=\left(\frac{\alpha\Phi}{r^{2}},\mp\alpha,\vec{0}\right),$

### v3

@@ 3.1 Volume complexity for the anisotropic case
Where, length scale of noncommutativity is $a=s^{1/4}\sqrt{\theta}$, $s$ is (dimensionless) t’Hooft coupling, $\sqrt{\theta}$ is the noncommutativity parameter and $B_{xy}$ is the only non-zero component of the NS-NS form background. The volume complexity can be found by taking a space like hypersurface specified by the embedding $t=t(u)$ in the string frame as,
@@
Here again we re-express the final result by replacing the bulk UV cutoff in terms of the boundary lattice spacing $\epsilon$ via $u_{b}=1/\epsilon$) and central charge $c$.
$\mathcal{C}_{V}(t_{*})$ $=\frac{1}{3G_{N}^{(5)}}\,\left(\frac{W}{\epsilon}\right)^{3}=\frac{8c}{3\pi}\left(\frac{W}{\epsilon}\right)^{3}$ (71)
@@ 3.2 Subregion Volume complexity (CV) for the anisotropic case
To compute subregion complexity we focus our attention on the portion of the maximal volume slice which is contained within the Ryu-Takayanagi (RT) surface, and which is also homologous to the boundary subregion of interest. We take the boundary subregion to be the infinitely long strip, $\frac{l}{2}<x<\frac{l}{2},\ -\frac{W}{2}<y,z<\frac{W}{2}$ in the limit $W\to\infty$ with $l$ fixed. In this case the pullback metric on the bounding codimension 2 (RT) surface is
@@ 3.2.1 Near AdS/Weakly noncommutative regime
$\mathcal{C}_{V}$ $=\frac{W^{2}}{G_{N}^{(5)}}\left(\frac{l}{3\epsilon^{3}}-\frac{4\pi^{3/2}\Gamma\left(\frac{2}{3}\right)}{9l^{2}\Gamma\left(\frac{1}{6}\right)}+a^{4}\left(\frac{128\pi^{9/2}\epsilon^{3}\Gamma\left(\frac{2}{3}\right)^{9}}{9l^{9}\Gamma\left(\frac{1}{6}\right)^{9}}-\frac{8\pi^{3/2}\Gamma\left(\frac{2}{3}\right)^{3}}{9l^{3}\epsilon^{3}\Gamma\left(\frac{1}{6}\right)^{3}}\right)+O(a^{8})\right)$ (86)
@@ 3.2.2 Far AdS/extreme noncommutative regime
$\mathcal{C}_{V}$ $=\frac{1}{G_{N}^{(5)}}\left(\frac{W^{2}l}{3\epsilon^{3}}+\frac{W^{2}l^{4}}{a^{6}}\frac{\sqrt{3}\Gamma\left(\frac{5}{6}\right)^{6}}{\sqrt[3]{2}\pi^{3}\Gamma\left(\frac{4}{3}\right)^{3}}-\frac{W^{2}l^{3}}{a^{4}\epsilon}\frac{\Gamma\left(-\frac{1}{6}\right)^{4}}{32\sqrt[3]{2}\sqrt{3}\pi^{3}\Gamma\left(\frac{1}{3}\right)}\right)$ (88)
@@ 3.3 Action complexity (CA) for the anisotropic case (10 D)
    Like we mentioned in the previous section, it turns out that we need to incorporate the appropriate topological boundary term Kurlyand:2022vzv ; Apolo:2025wcl to get correct action leading to the gravitational background with a negative cosmological constant.
@@
    $S_{\partial\mathcal{V}}$ $=\sum_{j=\pm}\frac{2}{16\pi G_{N}}\int d^{3}x\int ds_{j}\,\sqrt{\gamma_{j}}\,\kappa_{j}.$ (96)
    Here, the affine parameter $s$ along the null generators increases in the future direction, and $\gamma$ represents the induced metric on constant-$s$ slices. On the section of constant $s_{\pm}=\mp u$,
@@
    where the ++ and - indices denote the upper and lower null boundaries, respectively. The affine parameters are chosen such that $s_{\pm}=\mp u$, leading to the following null normal vectors
@@
    In the perturbative limit where $a\ll 1/u_{b}$, this expression receives vanishing contribution up to $\mathcal{O}(a^{4})$ in non-commutative coupling. This is consistent with the the expected AdS result.
    In contrast, the strongly non-commutative regime (where $a\gg 1/u_{b}$) contribution is given by
    $S_{\partial\mathcal{V}}$ $=\frac{L^{3}u_{b}^{3}}{16\pi G_{N}}\left(-\frac{4}{3}\alpha\right)+O\left(\frac{1}{a^{3}}\right).$ (98)
  • LMPS counterterm prescription: As noted in Lehner:2016vdi , the gravitational action depends on the choice of the affine parameter $s$ on null boundaries. To preserve reparametrization invariance, the following counterterm must be included for each null boundary
    $\sum_{\pm}\Delta S_{i}$ $=-2\int d{A_{i}}\,ds_{i}\,{\Theta_{i}}\ln|l_{ct}{\Theta_{i}}|,\hskip 85.35826pt\left({\Theta_{i}}\equiv\frac{1}{\sqrt{{\gamma}}}\frac{\partial\sqrt{{\gamma}}}{\partial s_{i}}\right).$
    Here, $\gamma$} denotes the induced metric on constant-$s$ cross-sections, which form codimension-2 surfaces. On slices with constant $s_{\pm}=\mp r$, the integration measure is as derived earlier ($\sqrt{\gamma}=u^{3}h(u)$). After accounting for sign conventions ($\Theta_{+}=-\Theta_{-}$), and to the causal structure, the total contribution from both surfaces turns out to be
    $\Delta S$ $=\frac{4}{16\pi G_{N}}\int d{A}\,\int_{0}^{\infty}ds_{-}\,{\Theta}_{-}\ln|{l_{ct}\Theta}_{-}|,$ (99)
@@
$\mathcal{C}_{\mathcal{A}}$ $=\frac{L^{3}}{16\pi^{2}G_{N}\epsilon^{3}}\left(4\ln\left(\frac{3l_{ct}}{\alpha}\right)\right)+O(a^{4}).$ (105)
Where in the final expression, we have conventionally expressed UV cutoff $u_{b}$ in terms of boundary lattice spacing $\epsilon$ via $u_{b}=1/\epsilon$ .
@@
$\mathcal{C}_{\mathcal{A}}$ $={\frac{L^{3}}{16\pi^{2}G_{N}\epsilon^{3}}\left(-2\ln(\alpha a)+\frac{20-4\alpha}{3}\right)+O\left({a^{-1}}\right).}$ (106)
The arguments of logarithm have been rendered dimensionless by rewriting them in the units of AdS radius. It can be seen that for the anisotropic case of noncommutative field theory, the action complexity is positive only in the range $\ln(\alpha a)<\frac{10-2\alpha}{3}$.This feature of action complexity is parallels with the action complexity of the DDSYM theory. However, unlike the DDSYM theory, but instead just like the local SYM field theory, the UV divergence scales with volume $L^{3}$! We defer our speculations about this counterintuitive result till Sec. 5.
@@ 4 Noncommutative SYM theory: Isotropic case
$ds_{E}^{2}=l^{2}\,{u^{2}}\,\Phi\,\left(\Phi^{-2}\left(dx_{0}^{2}+dx_{1}^{2}+dx_{2}^{2}+dx_{3}^{2}\right)+u^{-4}du^{2}+u^{-2}d\Omega^{2}_{5}\right)$
$\Phi$ $=(1+a^{4}u^{4})^{1/2},\quad F_{0123u}=i4l^{4}u^{3}\Phi^{-4},$
$B_{01}=B$ $_{23}=\sqrt{g_{s}}a^{2}u^{4}\Phi^{-2},\quad A_{01}=A_{23}=-i\frac{a^{2}l^{2}}{\sqrt{g_{s}}}u^{4}\Phi^{-2},\quad\chi=i\frac{a^{4}u^{4}}{g_{s}}$ (107)
@@
$ds^{2}$ $=l^{2}u^{-2}\Phi^{8/3}du^{2}+l^{2}u^{2}\Phi^{2/3}(-dt^{2}+dx_{1}^{2}+dx_{2}^{2}+dx_{3}^{2})$ (108)
where $\Phi=(1+a^{4}u^{4})^{1/2}$. We will work in spherical polar coordinates ($\rho,\theta,\phi$) instead of the cartesian $x_{1},x_{2},x_{3}$ and follow a similar procedure as above to compute the volume complexity.
@@ 4.1 Volume complexity of the isotropic NCSYM
$V$ $=\int du\ d\rho\ d\theta\ d\phi\ l^{4}u^{4}\Phi^{4/3}\rho^{2}\sin\theta\sqrt{\frac{\Phi^{2}}{u^{4}}-t^{\prime 2}(u)}=\frac{4\pi l^{4}R^{3}}{3}\int_{0}^{1/\epsilon}du\ u^{2}\Phi^{7/3}$ (109)
@@ 4.2 Action complexity (CA) for the isotropic case (5D)
$\mathcal{R}$ $=-\frac{20\left(3a^{8}u^{8}+10a^{4}u^{4}+3\right)}{3\left(a^{4}u^{4}+1\right)^{10/3}}.$
@@
$dt^{2}$ $=\frac{du^{2}}{u^{4}}\Phi^{2},$
$\int_{t_{-}}^{t_{+}}dt$ $=2\int^{u_{b}}_{u}du^{\prime}\,\frac{\Phi(u^{\prime})}{u^{\prime 2}}.$
@@
  • Computation of bulk action: The bulk action is evaluated over WdW patch with integration measure $\sqrt{-g}=u^{3}\Phi^{\frac{8}{3}}$ and gives
@@
    $=\frac{1}{16\pi G_{N}}\int d^{3}x\int_{0}^{u_{b}}du\int_{t_{-}}^{t_{+}}dt\,\,u^{3}\left(a^{4}u^{4}+1\right)^{4/3}\left(\frac{8\left(2-5\left(a^{4}u^{4}+1\right)^{2}\right)}{3\left(a^{4}u^{4}+1\right)^{10/3}}\right).$
    Working perturbatively, prior to capturing the non-commutative answer, we always first verify the AdS side of the story. Evaluating the bulk action in the perturbative limit when, the UV cutoff scale dominates the non-commuative coupling i.e. $a\ll 1/u_{b}$, reproduces the expected result for the known case of pure AdS
    $S_{bulk}$ $=\frac{V_{3}u_{b}^{3}}{16\pi G_{N}}\left(-\frac{4}{3}+O(a^{4})\right).$ (113)
    Once, we have the correct match with the expected scaling of the known AdS result, we next evaluate the non-commutative answer in the perturbative limit when the non-commutative coupling dominates the UV cutoff scale or in other words, $a\gg 1/u_{b}$. We obtain,
    $S_{bulk}$ $=\frac{V_{3}}{16\pi G_{N}}\left(-\frac{2}{3}a^{2}u_{b}^{5}+O(a^{-2})\right).$ (114)
@@
    $S_{boundary}$ $=\sum_{j=\pm}\frac{2}{16\pi G_{N}}\int d^{3}x\int ds_{j}\,\sqrt{\gamma_{j}}\,\kappa_{j}.$ (115)
    Here, the parameters $s$ on null generators are chosen to be increasing in the future direction, $\gamma$ is the induced metric on the cross-sections of constant $s$. Moreover, $n^{\mu}\nabla_{\mu}n^{\nu}=\kappa n^{\nu}$, where, $n$ is the null normal to the surface. Boundaries of WdW are given by Null surfaces $\mathcal{N}_{\pm}$
    $(t-T)$ $=\pm\int^{u_{b}}_{u}du^{\prime}\,\frac{\Phi(u^{\prime})}{u^{\prime 2}}.$
@@
    $t$ $=T\pm\int^{u_{b}}_{u}du^{\prime}\,\frac{\Phi(u^{\prime})}{u^{\prime 2}},$
    where, +/+/- labels the upper/lower null boundaries, respectively. The parameters on null generators are chosen to be increasing in the future direction and are taken to be $s_{\pm}=\mp u$. With this parametrization, the null normal vectors takes the following form
    $n_{\pm}^{\mu}=\left(\frac{\alpha\Phi(u)}{u^{2}},\mp\alpha,\vec{0}\right),$
@@
    $\kappa_{\pm}$ $=\mp\frac{10\alpha a^{4}u^{3}}{3\left(a^{4}u^{4}+1\right)}.$
@@
    $S_{boundary}$ $=\frac{1}{16\pi G_{N}}\frac{40\alpha V_{3}a^{4}}{3}\int_{0}^{u_{b}}du\,\frac{u^{6}}{\left(a^{4}u^{4}+1\right)},$
    where, on the sections of constant $s_{\pm}$, the induced metric is
    $d\gamma^{2}$ $=u^{2}(dx^{2}+dy^{2}+dz^{2})\Phi^{2/3},$
    giving the integral measure of $\sqrt{\gamma}=u^{3}\Phi(u)$. Working perturbatively, we first verify the AdS side of the story by evaluating the bulk action in the perturbative limit when, $a\ll 1/u_{b}$
    $S_{boundary}$ $=\frac{V_{3}}{16\pi G_{N}}a^{4}\alpha u_{b}^{7}\left(\frac{40}{21}+O(a^{4})\right),$ (116)
    this correctly reproduces the expected result for the known case of pure AdS to the zeroth order in coupling $a$. Now looking ahead and working out the corresponding answer when the non-commutative coupling dominates the UV cutoff scale $a\gg 1/u_{b}$. We obtain,
    $S_{boundary}$ $=\frac{V_{3}}{16\pi G_{N}}\left(\frac{40}{9}\alpha u_{b}^{3}+O\left(\frac{1}{a^{4}}\right)\right).$ (117)
@@
  • LMPS counter-term prescription It has been pointed out in Lehner:2016vdi that the gravitational action depends upon the choice of the parametrization $s$ on the null boundary components. To restore the coordinate independence of the gravitational action, it is required to include the following counterterm component in the action, for each of the null boundaries.
    $\Delta S$ $=\sum_{j=\pm}\frac{-2}{16\pi G_{N}}\int d{A}_{j}\,ds_{j}\,{\Theta}_{j}\ln|l_{ct}{\Theta}_{j}|,\hskip 85.35826pt\text{where,}\,\,{\Theta}_{\pm}\equiv\frac{1}{\sqrt{{\gamma}}}\frac{\partial\sqrt{{\gamma}}}{\partial s_{\pm}}.$ (118)
    Here, $\gamma$} is the induced metric on the constant $s$ sections. Since, metric is defined on the null section, it is a codim-2 surface. And on the section of constant $s_{\pm}=\mp u$, the integration measure is the same as obtained earlier.
@@
    $\text{ }\,\,\Theta$ $=\frac{5a^{4}u^{4}+3}{u\left(a^{4}u^{4}+1\right)}.$
@@
    $\Delta S$ $=\frac{4V_{3}}{16\pi G_{N}}\,\int^{\infty}_{0}du\,\sqrt{\gamma}\,{\Theta}\ln|l_{ct}{\Theta}|,$
    $=\frac{4V_{3}}{16\pi G_{N}}\,\int_{0}^{u_{b}}du\,u^{2}\frac{5a^{4}u^{4}+3}{\left(a^{4}u^{4}+1\right)^{1/2}}\ln\left|\frac{l_{ct}(5a^{4}u^{4}+3)}{u\left(a^{4}u^{4}+1\right)}\right|.$
@@
    $\Delta S$ $=\frac{V_{3}u_{b}^{3}}{16\pi G_{N}}\left[\frac{4}{3}+4\ln\left(\frac{3l_{ct}}{u_{b}}\right)\right]+O(a^{4}).$
@@
    $\Delta S$ $=\frac{V_{3}}{16\pi G_{N}}\left(2a^{2}u_{b}^{5}\left(-\ln\left(\frac{u_{b}}{l_{ct}}\right)+\ln\left(5\right)+1\right)\right)+O(a^{-2}).$ (119)
@@
    $S_{\mathcal{X}}$ $=\frac{-2}{16\pi G_{N}}\int d^{3}x\,\sqrt{\gamma}\ln\left|\frac{n_{+}\cdot n_{-}}{2}\right|\Bigg|_{u=u_{b}}.$
@@
    $d\gamma^{2}$ $=u^{2}(dx^{2}+dy^{2}+dz^{2})\Phi^{2/3},$
    With the integral measure $\sqrt{\gamma}=u^{3}\Phi(u)$. Given that the various normals have the following components,
    $n^{\mu}_{\pm}$ $=\left(\frac{\alpha\Phi}{u^{2}},\mp\alpha,\vec{0}\right),$